Given the input clauses S and the order a > b > c, produce a resolution proof (generating the empty clause) using only input resolution (unit resolution, or ordered resolution)
S:
a v ~p v ~c
b
a v ~b v c
~a v ~b
Answer #1

Solution :

c1 : a v ~b v ~c

c2 : b

c3 : a v ~b v c

c4 : ~a v ~b

Resolution Proof :

c2 + c4 = ~a (= c5)

c2 + c3 = a v c (= c6)

c5 + c6 = c (= c7)

c1 + c7 = a v ~b (= c8)

c5 + c8 = ~b (=c9)

c2 + c9 = FALSE .........empty clause